mike weights 90 pounds. On a web site, he calculated that he would weigh about 15 pounds on the Moon. Write an equation relating pounds on Earth, e, to pounds on the Moon, m. About how many pounds would Mike's dog weigh on the Moon if he weighs 54 pounds on Earth?
Answer by stanbon(75887)

mike weights 90 pounds. On a web site, he calculated that he would weigh about 15 pounds on the Moon. Write an equation relating pounds on Earth, e, to pounds on the Moon, m. About how many pounds would Mike's dog weigh on the Moon if he weighs 54 pounds on Earth?
------
m = (15/90)e
------
m(54) = (15/90)54 = (1/6)54 = 9 lbs (dog's wt on moon)
-------------
Cheers,
Stan H.
---------------